OpenStudy (anonymous):

HOW WE SOLVE THIS \[log_{2}(x+3)+log_{3}(x+2)=1\]

OpenStudy (anonymous):

I know the answer is -1

OpenStudy (anonymous):

I am sure @bhaskarbabu will do this..

OpenStudy (anonymous):

ya its -1 its little bit lengthy to explain just wait for a while

OpenStudy (anonymous):

loga + logb = logab First use this...

OpenStudy (anonymous):

\[\log_{y} x = \ln x /\] Then use this

OpenStudy (anonymous):

ln y

OpenStudy (anonymous):

So, we get (x+3)(x+2) = 2 You get a quadratic equation...Solve

hartnn (hartnn):

can't do that. tthe base for 2nd is 3

OpenStudy (anonymous):

You get -1 and -4

OpenStudy (anonymous):

But -4 is not valid as it does not follow rules of log

OpenStudy (anonymous):

Oh sorry..I saw it as 3..

OpenStudy (anonymous):

log(x+3) 2 = log3 3 - log(x+2) 3 log(x+3) 2 = log(3/x+2) 3 ln(x+3) ln2 =ln(3/x+2) ln3 ln3 * ln(x+3) = ln(2) * ln(3/x+2) hence x+3=2 and 3/x+2 =3

OpenStudy (anonymous):

hence x=-1

hartnn (hartnn):

can't understand that :(

ganeshie8 (ganeshie8):

\(log_2 (x+3) = log_3 3 - log_3(x+2) \\ log_2 (x+3) = log_3 (3/x+2)\)

ganeshie8 (ganeshie8):

till here, i understood ur work @bhaskarbabu

OpenStudy (anonymous):

He equaled the variable terms on side to the constant terms on the other side...Its a good way actually..

OpenStudy (anonymous):

@ganeshie8 after that he made all the bases to e..that is he changed them into natural logarithms..

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

next u converted everything to natural log, but hw u conclude x from there ?

OpenStudy (anonymous):

now cross multiply

OpenStudy (anonymous):

\[\log_{y} x = \ln x / \ln y\]

OpenStudy (anonymous):

@bhaskarbabu I get it. Excellent work. wow!!!!!!!

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

wow !! i see it too :) amazing work !!!!